1. A toy balloon has an internal pressure of 1.05 atm and avolume of 5.0...

A toy balloon has an internal pressure of \(1.05\) atm and a volume of \(5.0 \mathrm{~L}\). If the temperature where the balloon is released is \(20^{\circ} \mathrm{C}\), what will happen to the volume when the balloon rises to an altitude where the pressure is \(0.65\) atm and the temperature is \(-15^{\circ} \mathrm{C} ?\)
Answer #1

Given:

Pi = 1.05 atm

Pf = 0.65 atm

Vi = 5.0 L

Ti = 200.0 oC

= (200.0+273) K

= 473 K

Tf = -150.0 oC

= (-150.0+273) K

= 123 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(1.05 atm*5 L)/(473.0 K) = (0.65 atm*Vf)/(123.0 K)

Vf = 2.1003 L

Answer: 2.10 L